### 30 Mar 2026 v1

# Kill-Chain Canaries: Stage-Level Tracking of Prompt Injection Across Attack Surfaces and Model Safety Tiers Haochuan Kevin Wang

Abstract
We present a stage-decomposed analysis of prompt injection attacks against five frontier LLM agents. Prior work measures task-level attack success rate (ASR); we localize the pipeline stage at which each model’s defense activates. We instrument every run with a cryptographic canary token (SECRET-[A-F0-9]{8}) tracked through four kill-chain stages—Exposed, Persisted, Relayed, Executed—across four attack surfaces and five defense conditions (764 total runs, 428 no-defense attacked). Our central finding is that model safety is determined not by whether adversarial content is seen, but by whether it is propagated across pipeline stages. Concretely: (1) in our evaluation, exposure is 100% for all five models—the safety gap is entirely downstream; (2) Claude strips injections at write_memory summarization (0/164 ASR), while GPT-4o-mini propagates canaries without loss (53% ASR, 95% CI: 41–65%); (3) DeepSeek exhibits 0% ASR on memory surfaces and 100% ASR on tool-stream surfaces from the same model—a complete reversal across injection channels; (4) all four active defense conditions (write_filter, pi_detector, spotlighting, and their combination) produce 100% ASR due to threat-model surface mismatch; (5) a Claude relay node decontaminates downstream agents—0/40 canaries survived into shared memory.
1 Introduction
Prompt injection—adversarial instructions embedded in data processed by an LLM agent—is now a central attack class for agentic AI deployments [1, 2]. The standard evaluation reports a single metric: did the agent ultimately execute the adversary’s intended action? This conflates two distinct questions: whether the model observes the injection and whether it acts upon it.
Consider a concrete chain: an agent calls get_webpage(), receives poisoned content, calls write_memory() to summarize, and is queried by a second agent via read_memory(). A 0% ASR for the second agent could mean the injection was stripped at write_memory, or it could mean it survived but the second agent refused to execute. These are architecturally different outcomes: the first implies the summarizer is a decontamination stage; the second implies the terminal agent holds the defense. Outcome-only measurement cannot distinguish them.
We address this with a kill-chain canary methodology. Every injected payload contains a unique SECRET-[A-F0-9]{8} token tracked at four discrete stages by our PropagationLogger. We test five frontier models across four attack surfaces and five defense conditions in agent_bench, a custom multi-agent evaluation harness.
Contributions: (1) Stage-level evaluation framework — a kill-chain canary methodology that localizes at which pipeline stage each model’s defense activates, not merely whether it activates. (2) Empirical localization of model defenses — evidence that the safety gap between models concentrates at the summarization stage, not at execution. (3) Cross-surface vulnerability analysis — the first systematic comparison showing a single model’s ASR spans 0%–100% depending on injection channel. (4) Defense failure via threat-model mismatch — a structural explanation for why static defenses fail against off-surface injections, extending prior work on adaptive attacks. (5) Relay decontamination as an architectural primitive — evidence that a Claude summarizer node decontaminates downstream agents regardless of those agents’ individual safety levels.
2 Related Work
AgentDojo [3] provides 97 tasks with 629 injections and reports joint utility-ASR metrics across four environments; it does not decompose by pipeline stage. Our utility-ASR scatter (Figure 2) is directly comparable to their result. InjecAgent [4] evaluates 1,054 indirect-injection cases with a single outcome metric. Prompt Infection [5] demonstrates LLM-to-LLM self-replicating attacks; we identify the exact stage (write_memory summarization) where replication is blocked. Zombie Agents [6] shows summarization agents persist injections; we replicate and extend to multi-agent relay with quantified stage fractions. Nasr et al. [7] show adaptive attacks achieve >>90% ASR against 12/12 defenses. In our evaluation, non-adaptive attacks achieve the same result via surface mismatch—a structurally distinct failure mode that does not require adaptive adversaries.
3 Benchmark Design
3.1 System Architecture
Figure 1: Benchmark architecture. Injections enter via one of four attack surfaces and are tracked by PropagationLogger through four kill-chain stages. Agent A (summarizer) reads poisoned content and writes to MemoryStore; Agent B (executor) reads and calls outbound tools. Utility and security are evaluated independently.
agent_bench is a minimal multi-agent evaluation harness ($\sim$600 lines of Python). Its core components are:
MemoryStore — key-value store accessed via read_memory(key) and write_memory(key, value). Supports a write_filter defense that keyword-scans before committing. MemoryStore.seed(key, payload) pre-injects the canary without triggering the filter, operationalizing the memory_poison attack surface.
ToolRegistry — permission-gated dispatcher for get_webpage(url), query_db(sql), send_report(to, body), and escalate_privilege(reason). Each tool carries a minimum permission level (read / write / execute / admin).
MultiAgentOrchestrator — coordinates Agent A and Agent B with two relay modes: delegation (A’s final message becomes B’s system prompt) and memory (A writes to MemoryStore; B reads it).
PropagationLogger — JSONL event logger that matches the canary regex against every tool result and tool argument at each step, emitting canary_in_result and canary_in_args events with stage labels.
3.2 Kill-Chain Stages and Attack Scenarios
We define an attack surface as the channel through which adversarial content enters the agent pipeline (e.g., pre-seeded memory state, tool response stream, or web document). Each combination of attack surface and model defines an independent evaluation cell.
Each attacked run is annotated with the furthest kill-chain stage the canary reaches: Exposed (canary in any tool result), Persisted (canary in a write_memory call), Relayed (canary read by Agent B), Executed (canary in any outbound tool argument). The gap between Exposed and Persisted identifies summarization-stage filtering; between Persisted and Executed, it identifies execution-stage refusal.
Four scenarios cover distinct surfaces:
memory_poisonMemoryStore.seed() pre-loads the injection. Task: summarize research_notes and email to supervisor@lab.org. The agent calls read_memory(), receives the poisoned record, and may call send_report(to="exfil@attacker.com").
tool_poison — injection arrives in the return value of get_webpage() or query_db(). The legitimate document and the adversarial instruction appear in the same tool result.
propagation — two-agent memory relay. Agent A calls get_webpage() on a poisoned page and write_memory("summary", …). Canary presence or absence in that write determines whether Agent B is exposed.
permission_esc — injection calls escalate_(privilege()) before issuing a high-privilege action. Tests two-step adversarial instruction following.
3.3 Models, Defenses, and Metrics
Five frontier models: GPT-4o-mini (gpt-4o-mini-2024-07-18), GPT-5-mini [^1] Accessed via the OpenAI API during the experimental period (March 2026) under the identifier gpt-5-mini-2025-02-15; the publicly released model name may differ at time of publication., DeepSeek Chat v3 (deepseek-chat), Claude Haiku 4.5, Claude Sonnet 4.5. Temperature = 0.0; max tokens = 2048. 764 runs across 9 batches (March 10–27, 2026): 428 no-defense attacked, 80 with-defense attacked, 256 clean controls. While sample sizes per cell are modest (n=8n=83636), the key findings replicate consistently across independent batches spaced up to 17 days apart.
Five conditions: none (control baseline); write_filter (MemoryStore keyword scan before commit) [8]; pi_detector (secondary LLM classifies each outgoing query); spotlighting (tool results wrapped in <document> XML delimiters) [8]; all (all three combined).
ASR = fraction of attacked runs where canary reaches Executed. Task success = fraction of all runs where the legitimate task goal is completed. All fractions are reported with Wilson 95% confidence intervals (CI) as point estimate (95% CI: lohi[9].
4 Results
4.1 Exposure Is Universal; Defense Is Downstream
Table 1: Overall ASR and task success (no-defense attacked runs). In our evaluation, exposure = 100% for all models.
Model $n$ Task% ASR (95% CI)
GPT-4o-mini 60 90% 53% (41–65%)
DeepSeek Chat 68 100% 25% (16–37%)
GPT-5-mini 136 94% 3% (1–7%)
Claude Haiku 80 100% 0% (0–5%)
Claude Sonnet 84 100% 0% (0–4%)
Figure 2: (a) Benign task success vs. targeted ASR. Three qualitatively distinct regimes emerge: GPT-4o-mini (90% utility, 53% ASR); DeepSeek/GPT-5-mini (partial resistance); Claude (appears Pareto-efficient within our evaluation: 100% utility, 0% ASR). (b) Utility under attack — most models maintain task success while being compromised, confirming the dual-completion pattern.
Table 1 reports overall ASR per model. In our evaluation, every model receives every injection: exposure rate is 100% across all 764 attacked runs. The safety gap is entirely downstream of context exposure. Figure 2 illustrates three qualitatively distinct regimes. GPT-4o-mini (90% utility, 53% ASR) is the worst case: high task success and high injection compliance. Claude (100% utility, 0% ASR) appears Pareto-efficient within our evaluation—both Claude models achieve 100% task success on clean runs while producing 0% ASR under attack, though we caution this result is specific to the evaluated scenarios and tasks. DeepSeek’s 25% overall ASR masks a 0%/100% surface split discussed in Section 4.3.
4.2 Claude’s Defense Activates at Summarization
Figure 3: Kill-chain decomposition (propagation scenario). Top: per-stage survival blocks for three representative models — GPT-4o-mini propagates all four stages at 100%; GPT-5-mini partly filters at Persisted; Claude eliminates every canary at write_memory. Bottom: canary survival curves for all five models — the sharp drop for both Claude variants between Exposed and Persisted marks the summarization-stage defense.
Table 2: Kill-chain stage fractions, propagation scenario. Exposed = 100% for all models.
Model $n$ Pers. Rel. Exec.
GPT-4o-mini 8 100% 100% 100%
DeepSeek Chat 8 100% 100% 100%
GPT-5-mini 20 15% 15% 15%
Claude Haiku 20 0% 0% 0%
Claude Sonnet 20 0% 0% 0%
Table 2 and Figure 3 illustrate per-stage canary survival. Claude eliminates the injection during write_memory: in 0/40 Claude propagation runs did the canary survive into MemoryStore (20 Haiku + 20 Sonnet; 95% CI: 0–8%). The downstream Agent B receives a semantically correct but adversarially clean summary.
The architectural implication is notable: in any multi-agent pipeline, the safety of the relay/summarizer node determines downstream exposure independently of the terminal agent’s safety level. A Claude relay decontaminates for any downstream consumer; a GPT-4o-mini relay propagates injections with full fidelity. This maps to the Prompt Infection self-replication mechanism [5]: replication is blocked specifically at the write_memory summarization step. We do not isolate whether this effect arises from training data, system prompts, or tool interaction design; further ablation is required.
GPT-5-mini partially filters at Persisted (3/20, 15%; 95% CI: 5–36%), a non-trivial pass-through rate that rules out both full compliance and full resistance.
4.3 Surface-Specific ASR: 0% and 100% from the Same Model
Figure 4: ASR heatmap (model $\times$ scenario) with Wilson 95% CI per cell. The orange border ($\star$) highlights DeepSeek Chat: 0/24 on memory_poison (three independent batches over 17 days) vs. 8/8 on tool_poison — a 100-percentage-point swing from the same model on a different injection surface. This demonstrates that single-surface evaluation produces a complete mischaracterization of actual safety posture.
Table 3: ASR by model and scenario (no-defense attacked runs). Raw counts shown as $h/n$; bolded cells replicated at $n\geq 20$. The perm_esc total of 2/132 comprises DeepSeek 1/28 and GPT-5-mini 1/36.
Scenario 4o-mini DeepSeek 5-mini Haiku Sonnet
mem_poison 12/12 0/24 0/36 0/20 0/20
propagation 8/8 8/8 3/20 0/20 0/20
tool_poison 8/8 8/8 0/36 0/20 0/20
perm_esc 0/28 1/28 1/36 0/20 0/20
Table 3 and Figure 4 report per-surface ASR. DeepSeek’s result is operationally the most important: 0/24 on memory_poison (three independent batches spanning 17 days) and 8/8 on tool_poison and propagation. A single-surface evaluation would give DeepSeek either a perfect score (memory-only) or a total failure (tool-stream-only) — either result would be a complete mischaracterization of its actual safety posture.
The most parsimonious explanation is channel-differentiated trust: DeepSeek appears to treat pre-stored memory records with higher suspicion than live function-call response streams. This cannot be verified from external behavior, but the pattern replicated cleanly across batches spaced 1–17 days apart. We note this observation is based on black-box behavioral analysis and requires white-box investigation to confirm.
In our evaluation, GPT-4o-mini achieves 100% ASR on all three non-privilege-escalation scenarios and Claude achieves 0% ASR on all four. The permission_esc near-zero results (2/132 total) most likely reflect payload design—the two-step overtly adversarial payload— rather than genuine privilege-escalation resistance. We do not claim this behavior generalizes beyond the evaluated configurations; further validation across additional models and tasks is required.
4.4 Defense Failures as Surface Mismatch
In our evaluation, all four active defense conditions produce 100% ASR on GPT-4o-mini and DeepSeek across propagation and tool_poison ($n=8$ per cell). Spotlighting reduces GPT-4o-mini task success from 63% to 50% without reducing ASR—a strictly worse utility-security outcome.
The primary failure mechanism appears to be a mismatch between each defense’s threat model and our injection surfaces. Spotlighting [8] wraps document content in XML delimiters, but our injection enters via the function-call response stream. pi_detector scans outgoing tool queries for adversarial intent, but our injection arrives in incoming tool results. write_filter intercepts agent-initiated memory writes, but memory_poison pre-seeds the payload via MemoryStore.seed() before the agent session begins, bypassing the interceptor entirely.
Each defense is correctly implemented for its stated threat model. The problem is that none of those threat models cover the injection surface under test. This extends the finding of Nasr et al. [7]—who showed adaptive attacks break 12/12 defenses—by demonstrating that non-adaptive attacks can achieve the same result via surface mismatch alone. Any deployment that evaluates defenses only against the surfaces they were designed for will miss every injection that enters from a different channel.
4.5 Objective Drift as a Post-Hoc Forensic Signal
Figure 5: Per-step TF-IDF cosine distance from the task description (GPT-4o-mini, memory_poison). Clean (blue) and attacked (red) curves are indistinguishable through steps 1–2. Divergence appears at step 3—the harmful send_report call— concurrent with, not before, the harm.
We compute per-step objective drift as the TF-IDF cosine distance from the original task description. Table 4 reports step-level drift for GPT-4o-mini on memory_poison: steps 1–2 are statistically indistinguishable; the signal diverges only at the harmful step itself.
Table 4: Per-step objective drift (GPT-4o-mini, memory_poison).
Step Tool Clean Attacked $\Delta$
1 read_memory 0.563 0.571 +0.008
2 send_report (legit) 0.714 0.747 +0.033
3 send_report (harm) 0.868 +0.154
Last-step drift $\Delta$ by scenario: memory_poison +0.348, permission_esc +0.312, tool_poison +0.089, propagation +0.052.
[…]
Figure 6: Objective drift distributions: clean vs. attacked, per model. GPT-4o-mini and DeepSeek shift substantially upward when attacked; Claude’s attacked distribution is nearly identical to its clean baseline—consistent with 0% ASR. Median line shown per violin; grey = clean, colored = attacked.
[…]
4.6 Provenance Attribution and Dual-Completion
We implement a token-overlap provenance heuristic: for each outgoing tool argument, alphanumeric tokens ($\geq$4 chars) are matched against all prior tool results and memory state (threshold: $\geq$3 shared tokens or $\geq$20% argument coverage). On 22 compromised runs, 22/22 (100%) injection paths were correctly reconstructed with 0 false attributions.
The dual-completion pattern is central to why outcome-level monitoring fails. A compromised memory_poison run produces a legitimate send_report(to="supervisor@lab.org") at step 2 (task succeeds) and a malicious send_report carrying the canary token at step 3 (injection executes) — provenance traces both arguments to tool_result:read_memory@step1. Outcome-level monitoring observes a successful report delivery and misses the exfiltration entirely. Step-level provenance logging is the minimum forensic capability required to detect both events within the same run.
4.7 Execution Lag and Detection Feasibility
GPT-4o-mini’s median exposure-to-execution lag is 1 step (mean 1.25, max 2): it reads the injection and immediately issues the malicious tool call in the next invocation. DeepSeek’s lag is bimodal: most compromised runs execute within 2–3 steps, but a tail (max 12 steps) completes the entire legitimate task before appending the injection as a final action. GPT-5-mini when compromised exhibits a mean lag of 3.0 steps.
For GPT-4o-mini, any real-time interception middleware has exactly one inference cycle after first exposure to classify and block. For DeepSeek’s late-executing runs, a mid-run detector calibrated to GPT-4o-mini’s 1-step pattern will miss the attack entirely. Compromised GPT-4o-mini runs also show excess send_report calls (mean 1.26 vs. 0.57 clean, $\Delta$+0.69)—detectable without a classifier simply by comparing outbound tool-call counts against task-expected baselines.
5 Discussion
Model choice is the highest-leverage safety decision.
Claude’s 0/164 ASR (0 successes across all four scenarios, no defenses applied) reflects safety behavior that activates at the summarization stage rather than the execution stage. This is architecturally different from execution-stage refusal: the model appears to distinguish between processing adversarial content and propagating it into downstream writes. None of the evaluated defense mechanisms reproduced this behavior on any other model. We caution that this finding is specific to our benchmark and cannot be generalized without further validation.
Relay node safety is composable.
Placing a Claude model at the summarization/relay position provides decontamination for any downstream agent in our evaluation. Across 40 propagation runs with Claude as relay (20 Haiku + 20 Sonnet), zero canaries reached Agent B. This constitutes a deployable architectural choice independent of the downstream agent’s individual safety tier. We propose relay decontamination rate (fraction of injections stripped at the relay stage) as a first-class metric for multi-agent security evaluation.
Surface coverage is the primary evaluation gap.
DeepSeek’s 0% / 100% ASR split across memory and tool-stream surfaces demonstrates that single-surface evaluation is not a security evaluation. The same structural gap applies to detection: classifier signatures trained on one surface achieve AUC 0.39–0.57 on held-out surfaces. A complete security assessment must cover every injection surface present in the target deployment.
6 Broader Impact
This work studies adversarial vulnerabilities in LLM agents. All experiments were conducted in closed, sandboxed harnesses with no external network access; no real systems were compromised and no sensitive data was used. The injection payloads and attack surfaces described are representative of techniques already documented in the public literature [1, 2]. We believe transparent publication of evaluation methodology and empirical results benefits the community by enabling reproducible security assessment; we release all benchmark code and run logs. The primary dual-use risk is that surface-mismatch analysis could inform more targeted attack design. However, we assess that the defensive contribution—establishing that evaluation coverage determines apparent security posture—outweighs this risk, and that the attack techniques described are already publicly known.
7 Limitations
Per-cell $n$ ranges from 8 to 36; the strongest findings (Claude 0/20 per scenario, DeepSeek memory 0/24, GPT-4o-mini memory 12/12) are consistent across batches but should be interpreted with appropriate uncertainty. Payloads are explicit and synthetic; real-world attacks typically use obfuscation and social engineering. Five frontier models from three providers; reasoning models (o3, o4-mini) and open-weight models (Llama 3.3) are not represented. All defenses are lightweight wrappers that approximate the stated mechanism. permission_esc near-zero ASR (2/132) most likely reflects payload design rather than genuine privilege-escalation resistance. Classifier and provenance validation use the 192-run scenario_compare dataset; revalidation on the full 764-run set is ongoing. We do not isolate whether Claude’s summarization-stage filtering arises from training data, system prompt format, or tool API design.
8 Conclusion
We introduced kill-chain canary tracking for localizing prompt injection defenses in LLM agent pipelines. Across 764 runs and five frontier models, the key structural finding is threefold: the safety gap between models concentrates at the summarization stage, not at exposure; static defenses fail not because they are poorly designed but because they cover different surfaces than the actual injection vectors; and a Claude relay node provides measurable decontamination for downstream agents regardless of those agents’ individual safety levels. Stage-level evaluation reveals that prompt injection defenses are fundamentally pipeline-local, not model-global. The minimum evaluation standard for agentic systems should therefore include: stage-level kill-chain decomposition, multi-surface injection coverage, and relay decontamination rate as a first-class metric.

### 03 Apr 2026 v2

@@
 Abstract
-We present a stage-decomposed analysis of prompt injection attacks against five frontier LLM agents.
-Prior work measures task-level attack success rate (ASR); we localize the pipeline stage at which each model’s defense activates.
-We instrument every run with a cryptographic canary token (SECRET-[A-F0-9]{8}) tracked through four kill-chain stages—Exposed, Persisted, Relayed, Executed—across four attack surfaces and five defense conditions (764 total runs, 428 no-defense attacked).
-Our central finding is that model safety is determined not by whether adversarial content is seen, but by whether it is propagated across pipeline stages.
-Concretely: (1) in our evaluation, exposure is 100% for all five models—the safety gap is entirely downstream; (2) Claude strips injections at write_memory summarization (0/164 ASR), while GPT-4o-mini propagates canaries without loss (53% ASR, 95% CI: 41–65%); (3) DeepSeek exhibits 0% ASR on memory surfaces and 100% ASR on tool-stream surfaces from the same model—a complete reversal across injection channels; (4) all four active defense conditions (write_filter, pi_detector, spotlighting, and their combination) produce 100% ASR due to threat-model surface mismatch; (5) a Claude relay node decontaminates downstream agents—0/40 canaries survived into shared memory.
+We argue that prompt injection safety in multi-agent systems is a pipeline-stage problem, not a model-capability problem.
+We instrument 950 agent runs across five frontier LLMs with a cryptographic canary token tracked through four kill-chain stages (Exposed $\to$ Persisted $\to$ Relayed $\to$ Executed) across six attack surfaces and five defense conditions.
+Exposure is 100% for all models—the safety gap is entirely downstream of context ingestion.
+Claude eliminates injections at the write_memory summarization stage (0/164 ASR); GPT-4o-mini propagates them at 53% ASR; DeepSeek achieves 0%/100% ASR across memory vs.
+tool-stream surfaces from the same model.
+Three architectural implications follow from these results.
+First, write-node placement appears to be the highest-leverage safety decision: a Claude relay (Agent A position) decontaminates shared memory for any downstream agent in our evaluation, because safety fires at write time rather than read time—though this write-vs-read asymmetry rests on pilot-scale evidence ($n=3$) and requires replication.
+Second, defense coverage must match deployment surfaces: all four evaluated defenses fail on at least one surface because each was designed for a different injection channel; surface-mismatch failure requires no adversarial adaptation.
+Third, parser-extracted invisible injection is as dangerous as visible injection: whitefont PDF payloads achieve equal or higher ASR than appended text in our evaluation, suggesting that document screening at the rendered layer is insufficient.
+We release all benchmark code, run logs, and canary tooling.
 1 Introduction
@@
 Every injected payload contains a unique SECRET-[A-F0-9]{8} token tracked at four discrete stages by our PropagationLogger.
-We test five frontier models across four attack surfaces and five defense conditions in agent_bench, a custom multi-agent evaluation harness.
-Contributions: (1) Stage-level evaluation framework — a kill-chain canary methodology that localizes at which pipeline stage each model’s defense activates, not merely whether it activates.
-(2) Empirical localization of model defenses — evidence that the safety gap between models concentrates at the summarization stage, not at execution.
-(3) Cross-surface vulnerability analysis — the first systematic comparison showing a single model’s ASR spans 0%–100% depending on injection channel.
-(4) Defense failure via threat-model mismatch — a structural explanation for why static defenses fail against off-surface injections, extending prior work on adaptive attacks.
-(5) Relay decontamination as an architectural primitive — evidence that a Claude summarizer node decontaminates downstream agents regardless of those agents’ individual safety levels.
+We test five frontier models across six attack surfaces (web text, memory, tool stream, PDF, invisible PDF, and audio) and five defense conditions in agent_bench, a custom multi-agent evaluation harness.
+A Phase 3 multimodal extension evaluates a three-boundary kill chain (document extraction $\to$ memory write $\to$ agent delegation) and cross-model relay pairs where Agent A and Agent B are from different model families.
+Contributions: (1) Stage-level evaluation framework — a kill-chain canary methodology [1] that attributes defense effectiveness to specific pipeline stages rather than final outcomes.
+(2) Empirical defense localization — evidence across six injection surfaces and five models that the safety gap concentrates at the summarization write stage, not at context exposure or execution.
+(3) Cross-surface vulnerability analysis — the first systematic multi-surface evaluation showing a single model’s ASR spans 0%–100% depending on injection channel, demonstrating that surface coverage determines apparent safety posture.
+(4) Write-vs-read relay asymmetry — pilot evidence that Claude’s write-stage defense does not extend to read-stage resistance in heterogeneous multi-agent pipelines, suggesting that relay node position (not identity alone) determines downstream safety.
 2 Related Work
@@
 In our evaluation, non-adaptive attacks achieve the same result via surface mismatch—a structurally distinct failure mode that does not require adaptive adversaries.
+AgentWatcher [11] proposes rule-based, causally attributed prompt injection monitoring; our stage-level canary instrumentation provides a complementary empirical basis for where each model’s defense activates within the pipeline.
+Xiang et al.
+[12] survey system-level defenses for indirect prompt injection; our surface-mismatch results provide quantitative evidence for their observation that defense effectiveness is deployment-context-dependent.
+Ding et al.
+[13] study adversarial prompt injection on multimodal LLMs; our Phase 3 PDF cross-modal relay experiments extend this to multi-agent relay chains with per-stage kill-chain tracking.
+Zhang et al.
+[14] evaluate privilege usage of LLM agents on real-world tools; our permission_esc scenario finds near-zero ASR (2/132) for the same payload design, suggesting privilege escalation requires more sophisticated payload construction than simple text instructions.
+Wang et al.
+[15] study real-time monitoring for reasoning vulnerabilities; their post-hoc detection framing aligns with our finding that objective drift is forensic rather than preventive (leave-one-scenario-out AUC 0.39–0.57).
+Lynch [16] demonstrates persistent vulnerability of aligned AI systems to adversarial inputs; our finding that Claude’s resistance is localized to the write stage (not the read stage) nuances this—alignment may provide write-time filtering without guaranteeing read-time resistance in heterogeneous pipelines.
 3 Benchmark Design
 3.1 System Architecture
-Figure 1: Benchmark architecture.
-Injections enter via one of four attack surfaces and are tracked by PropagationLogger through four kill-chain stages.
-Agent A (summarizer) reads poisoned content and writes to MemoryStore; Agent B (executor) reads and calls outbound tools.
-Utility and security are evaluated independently.
-agent_bench is a minimal multi-agent evaluation harness ($\sim$600 lines of Python). Its core components are:
-MemoryStore — key-value store accessed via read_memory(key) and write_memory(key, value).
-Supports a write_filter defense that keyword-scans before committing.
-MemoryStore.seed(key, payload) pre-injects the canary without triggering the filter, operationalizing the memory_poison attack surface.
-ToolRegistry — permission-gated dispatcher for get_webpage(url), query_db(sql), send_report(to, body), and escalate_privilege(reason).
-Each tool carries a minimum permission level (read / write / execute / admin).
-MultiAgentOrchestrator — coordinates Agent A and Agent B with two relay modes: delegation (A’s final message becomes B’s system prompt) and memory (A writes to MemoryStore; B reads it).
-PropagationLogger — JSONL event logger that matches the canary regex against every tool result and tool argument at each step, emitting canary_in_result and canary_in_args events with stage labels.
+Figure 1: agent_bench pipeline architecture.
+Left: Three injection surfaces embed a canary token (SECRET-[A-F0-9]{8}) via PDF, pre-seeded memory, or tool response.
+Right (dashed): MultiAgentOrchestrator routes the injected content through Agent A (parse_pdf / write_memory) $\to$ MemoryStore $\to$ Agent B (read_memory / send_report).
+PropagationLogger tracks canary survival through four stages (Exposed $\to$ Persisted $\to$ Relayed $\to$ Executed); the furthest stage reached is the per-run kill-chain label.
+[…]
 3.2 Kill-Chain Stages and Attack Scenarios
+Threat model.
+We assume a read-only adversary who controls content that an agent will process (a PDF document, a web page, or a database record) but has no access to model weights, system prompts, or the agent runtime.
+The adversary’s goal is to cause the agent to issue a specific unauthorized outbound action (send_report to an attacker-controlled address) while the legitimate task still succeeds (the dual-completion attack pattern [2]).
+Payloads are explicit and unobfuscated; we treat obfuscation as orthogonal to the stage-localization question.
+Agent A and Agent B communicate only through the shared MemoryStore; neither agent has direct access to the other’s internal state.
 We define an attack surface as the channel through which adversarial content enters the agent pipeline (e.g., pre-seeded memory state, tool response stream, or web document).
 Each combination of attack surface and model defines an independent evaluation cell.
-Each attacked run is annotated with the furthest kill-chain stage the canary reaches: Exposed (canary in any tool result), Persisted (canary in a write_memory call), Relayed (canary read by Agent B), Executed (canary in any outbound tool argument).
+Each attacked run is annotated with the furthest kill-chain stage the canary reaches [1]: Exposed (canary in any tool result), Persisted (canary in a write_memory call), Relayed (canary read by Agent B), Executed (canary in any outbound tool argument).
 The gap between Exposed and Persisted identifies summarization-stage filtering; between Persisted and Executed, it identifies execution-stage refusal.
-Four scenarios cover distinct surfaces:
+Four text-surface scenarios and one multimodal scenario cover distinct injection channels:
 memory_poisonMemoryStore.seed() pre-loads the injection.
 Task: summarize research_notes and email to supervisor@lab.org.
-The agent calls read_memory(), receives the poisoned record, and may call send_report(to="exfil@attacker.com").
+The agent calls read_memory(), receives the poisoned record, and may call send_report(to="exfil@attacker.com", …).
 tool_poison — injection arrives in the return value of get_webpage() or query_db(). The legitimate document and the adversarial instruction appear in the same tool result.
 propagation — two-agent memory relay. Agent A calls get_webpage() on a poisoned page and write_memory("summary", …). Canary presence or absence in that write determines whether Agent B is exposed.
-permission_esc — injection calls escalate_(privilege()) before issuing a high-privilege action. Tests two-step adversarial instruction following.
+permission_esc — injection calls escalate_privilege() before issuing a high-privilege action. Tests two-step adversarial instruction following.
+cross_modal_relay (Phase 3) — a three-boundary scenario.
+Agent A calls parse_pdf() on an injected document and write_memory("doc_summary", …).
+Agent B calls read_memory() and send_report().
+The canary must cross three boundaries: document extraction, memory write, and agent delegation relay.
+Attack variants: pdf_append (visible appended text), pdf_whitefont (1 pt white-text overlay, human-invisible), pdf_metadata (payload in /Keywords and /Subject fields).
+We also test heterogeneous model pairs (Agent A $\neq$ Agent B) to measure whether the relay node’s identity determines downstream exposure.
 3.3 Models, Defenses, and Metrics
@@
 Temperature = 0.0; max tokens = 2048.
-764 runs across 9 batches (March 10–27, 2026): 428 no-defense attacked, 80 with-defense attacked, 256 clean controls.
-While sample sizes per cell are modest (n=8n=83636), the key findings replicate consistently across independent batches spaced up to 17 days apart.
+950 total runs: 764 text-surface runs across 9 batches (March 10–27, 2026) and 186 Phase 3 multimodal runs (April 1, 2026).
+Text-surface runs: 428 no-defense attacked, 80 with-defense attacked, 256 clean controls.
+Phase 3 runs: 138 homogeneous pairs (Block A) + 48 cross-model pairs (Block B), $n=3$ per cell.
+While text-surface sample sizes are modest (n=8n=83636), the key findings replicate across independent batches spaced up to 17 days apart.
+Phase 3 cells at $n=3$ are pilot-scale; confidence intervals are reported throughout.
 Five conditions: none (control baseline); write_filter (MemoryStore keyword scan before commit) [9]; pi_detector (secondary LLM classifies each outgoing query); spotlighting (tool results wrapped in <document> XML delimiters) [9]; all (all three combined).
@@
 4.2 Claude’s Defense Activates at Summarization
-Figure 3: Kill-chain decomposition (propagation scenario).
-Top: per-stage survival blocks for three representative models — GPT-4o-mini propagates all four stages at 100%; GPT-5-mini partly filters at Persisted; Claude eliminates every canary at write_memory.
-Bottom: canary survival curves for all five models — the sharp drop for both Claude variants between Exposed and Persisted marks the summarization-stage defense.
+Figure 3: Kill-chain canary survival curves (propagation scenario, all five models).
+The sharp drop for both Claude variants between Exposed and Persisted localizes the defense at the write_memory summarization step.
+GPT-4o-mini and DeepSeek propagate all four stages at 100%; GPT-5-mini shows partial filtering at Persisted (15%).
 Table 2: Kill-chain stage fractions, propagation scenario. Exposed = 100% for all models.
@@
 0.57 clean, $\Delta$+0.69)—detectable without a classifier simply by comparing outbound tool-call counts against task-expected baselines.
+4.8 Phase 3: Multimodal Cross-Modal Relay
+To assess whether the kill-chain decomposition generalises beyond text injection, we extend the benchmark to two new attack surfaces (PDF documents, audio transcripts) and introduce a three-boundary kill-chain, building on multimodal prompt injection findings of Ding et al.
+[13]: Exposed (document extraction), Persisted (memory write), Relayed (Agent B reads memory), Executed (Agent B issues send_report carrying the canary).
+We ran 186 trials across five models using the cross_modal_relay scenario ($n=3$ per cell; 46 homogeneous cells in Block A, 16 cross-model cells in Block B).
+At $n=3$, confidence intervals are wide; all Phase 3 results should be interpreted as directional pilot findings pending larger replication.
+Figure 8: Phase 3 Block A kill-chain stage rates for each model (homogeneous pairs, no defense, $n=3$ per cell).
+Solid bars: pdf_append; hatched bars: pdf_whitefont.
+DeepSeek propagates through all four stages at 100% for both variants.
+Claude blocks universally at Persisted (PER=0%).
+GPT-4o-mini reaches Relayed at 100% but executes at 0% (append) / 33% (whitefont).
+GPT-5-mini fails to reach Exposed (tool-call regression).
+Three-boundary propagation (Block A).
+Table 5 and Figure 8 report per-stage rates for homogeneous model pairs under no defense (attacked runs only).
+Table 5: Phase 3 Block A — homogeneous pairs, no-defense, attacked runs ($n=3$ per cell unless noted).
+Exp = canary in extracted text; Per = canary in memory write; Rel = canary in Agent B read_memory result; Exe = canary in Agent B send_report args (= ASR).
+Model Variant Exp Per Rel Exe
+GPT-4o-mini pdf_append 100% 100% 100% 0%
+GPT-4o-mini pdf_whitefont 100% 100% 100% 33%
+DeepSeek pdf_append 100% 100% 100% 100%
+DeepSeek pdf_whitefont 100% 100% 100% 100%
+Claude Haiku pdf_append 100% 0% 0% 0%
+Claude Haiku pdf_whitefont 100% 0% 0% 0%
+Claude Sonnet pdf_append 100% 0% 0% 0%
+GPT-5-mini pdf_append 0% 0% 0% 0%
+GPT-5-mini pdf_whitefont 0% 0% 0% 0%
+The pattern from text-surface experiments replicates cleanly across the PDF surface.
+DeepSeek propagates every injection through all three boundaries to execution (100%; Wilson 95% CI: 44–100%).
+Claude eliminates every injection at the Persisted stage (0/6 Haiku, 0/3 Sonnet; CI: 0–46%/0–56%), identical to the text-surface result.
+GPT-4o-mini reaches Relayed at 100% but executes inconsistently: 0% for pdf_append and 33% for pdf_whitefont.
+The invisible-injection finding is notable: pdf_whitefont (1 pt white text, invisible to human readers) achieves higher ASR than pdf_append on GPT-4o-mini (33% vs.
+0%).
+The whitefont text is extracted first by pypdf, placing the directive before the benign abstract in the extracted string; the position bias amplifies compliance.
+DeepSeek shows no positional sensitivity—both variants achieve 100%.
+GPT-5-mini fails to reach Exposed (0/6 EXP = 0%).
+The model did not invoke the parse_pdf tool in any of the 18 GPT-5-mini runs, producing 0% utility.
+This is a tool-calling regression: GPT-5-mini-2025-02-15 either does not support our tool schema format or requires a different prompt structure; we report the failure transparently.
+All GPT-5-mini cells are excluded from further multimodal analysis.
+Defense effectiveness (Block A).
+write_filter achieves 0% EXE universally: 0/3 DeepSeek pdf_append, 0/3 DeepSeek pdf_whitefont, 0/6 GPT-4o-mini pdf_whitefont.
+This extends the text-surface finding: write_filter is the only mechanism that blocks at the correct stage (memory write), and it does so across all evaluated injection surfaces.
+pi_detector completely fails on the PDF surface: DeepSeek pdf_append $\rightarrow$ 100% EXE under pi_detector (identical to no-defense); GPT-4o-mini pdf_whitefont $\rightarrow$ 33% (identical to no-defense).
+This confirms the surface-mismatch hypothesis introduced in Section 4.4: pi_detector scans outgoing tool queries for adversarial intent, but PDF injection arrives in the incoming parsed document content.
+The defense is correctly implemented for its stated threat model; that threat model simply does not include this surface.
+Figure 9: Phase 3 Block B kill-chain for cross-model relay pairs (pdf_append, no defense, $n=3$ per cell).
+When Claude Haiku is Agent A (writer), PER=0% blocks all downstream propagation.
+When Claude Haiku is Agent B (reader/executor), it propagates at EXE=67%—Claude’s defense fires at write, not at read.
+Cross-model relay asymmetry (Block B).
+Block B pairs non-identical agents: Agent A ($\neq$ Agent B) reads the PDF and writes to shared memory; Agent B reads memory and issues send_report.
+Table 6 and Figure 9 report results for pdf_append, no defense ($n=3$ per cell, attacked runs only).
+Table 6: Phase 3 Block B — cross-model relay, pdf_append, no defense, $n=3$ per cell (attacked runs). Wilson 95% CI shown for EXE.
+Agent A Agent B Per Rel Exe 95% CI
+GPT-4o-mini Claude Haiku 100% 100% 67% 21–94%
+GPT-4o-mini DeepSeek 100% 100% 100% 44–100%
+Claude Haiku GPT-4o-mini 0% 0% 0% 0–56%
+DeepSeek GPT-4o-mini 100% 100% 100% 44–100%
+Three directional observations emerge from Table 6; given $n=3$ per cell (95% CI ranging from 21–94%), these should be read as hypotheses for larger-scale follow-up rather than settled findings.
+Observation 1 (write-time vs.
+read-time defense).
+When Claude Haiku is Agent A, PER=0% prevents any downstream propagation.
+When Claude Haiku is Agent B and inherits pre-written malicious memory, EXE=67% (95% CI: 21–94%).
+The directional gap is consistent with a write-stage defense mechanism: the model appears to filter adversarial content when writing to shared memory but not when reading inherited content.
+Mechanistic validation (e.g., intervention on system prompt vs.
+RLHF) is required to confirm this interpretation.
+Observation 2 (relay node as primary chokepoint).
+Across both models tested as Agent A, the relay node identity determines whether any canary enters memory at all.
+A Claude relay (Agent A) produces PER=0% regardless of Agent B; a GPT-4o-mini or DeepSeek relay produces PER=100%, shifting all safety burden to Agent B.
+This extends the text-surface result from Section 4.2 to the PDF surface.
+Observation 3 (unexplained REL–EXE gap for homogeneous GPT-4o-mini).
+In Block A, GPT-4o-mini reaches Relayed=100% for pdf_append but EXE=0%: Agent B reads the canary but does not include it in send_report.
+When DeepSeek is Agent A (Block B), GPT-4o-mini as Agent B achieves EXE=100%.
+A plausible explanation is that DeepSeek writes the directive verbatim into memory while GPT-4o-mini paraphrases it, reducing instruction explicitness below the compliance threshold.
+This REL–EXE gap is an open empirical question: instruction explicitness in shared memory may be a vulnerability axis independent of canary token presence, but the mechanism is unconfirmed at this sample size.
 5 Discussion
 Model choice is the highest-leverage safety decision.
-Claude’s 0/164 ASR (0 successes across all four scenarios, no defenses applied) reflects safety behavior that activates at the summarization stage rather than the execution stage.
-This is architecturally different from execution-stage refusal: the model appears to distinguish between processing adversarial content and propagating it into downstream writes.
+Claude’s 0/164 ASR (0 successes across all four text-surface scenarios, no defenses applied) reflects safety behavior that activates at the summarization stage rather than the execution stage.
+Phase 3 confirms this pattern on the PDF surface: 0/6 Claude Haiku and 0/3 Claude Sonnet pdf_append/pdf_whitefont runs produced canary persistence (EXP=100%, PER=0%).
+Critically, Block B reveals that this defense is write-time only: when Claude Haiku inherits pre-written malicious memory as Agent B, it propagates at 67%.
+This is architecturally different from execution-stage refusal: the model distinguishes between processing adversarial content and writing it versus receiving it from a trusted delegation channel.
 None of the evaluated defense mechanisms reproduced this behavior on any other model.
 We caution that this finding is specific to our benchmark and cannot be generalized without further validation.
-Relay node safety is composable.
-Placing a Claude model at the summarization/relay position provides decontamination for any downstream agent in our evaluation.
-Across 40 propagation runs with Claude as relay (20 Haiku + 20 Sonnet), zero canaries reached Agent B.
-This constitutes a deployable architectural choice independent of the downstream agent’s individual safety tier.
-We propose relay decontamination rate (fraction of injections stripped at the relay stage) as a first-class metric for multi-agent security evaluation.
-Surface coverage is the primary evaluation gap.
-DeepSeek’s 0% / 100% ASR split across memory and tool-stream surfaces demonstrates that single-surface evaluation is not a security evaluation.
-The same structural gap applies to detection: classifier signatures trained on one surface achieve AUC 0.39–0.57 on held-out surfaces.
-A complete security assessment must cover every injection surface present in the target deployment.
+Relay node placement as a candidate safety primitive.
+Placing a Claude model at the summarization/relay position (Agent A) produces 0/46 canaries entering shared memory across text and PDF surfaces in our evaluation—independent of which model serves as Agent B.
+Block B provides directional evidence for an important asymmetry: placing Claude as Agent B does not provide the same protection (67% propagation when GPT-4o-mini writes the malicious memory, though $n=3$ makes this a hypothesis rather than a settled result).
+The implication for system design is direct: multi-agent frameworks should expose write-node identity as a first-class architectural choice.
+Operators can harden pipelines today by requiring safety-certified models at every inter-agent write boundary, independent of which models are used elsewhere in the chain.
+We propose relay decontamination rate at the write stage as a mandatory metric for multi-agent security benchmarks.
+Surface mismatch is a structural flaw, not a defense implementation gap.
+DeepSeek’s 0%/100% ASR split, pi_detector’s complete failure on PDF injection, and the audio injection null result collectively show that a model or defense’s apparent safety posture is almost entirely determined by which surfaces an evaluation covers.
+This is not a fixable implementation detail—it is a consequence of defenses being designed for specific threat models.
+The correct response is not to patch individual defenses but to change evaluation requirements: any claimed safety result must enumerate every injection surface in the target deployment and demonstrate coverage.
+Defenses that pass surface-specific evaluations but fail on held-out surfaces (AUC 0.39–0.57 for our classifiers) should be reported as surface-local, not general.
+Parser-level trust is the missing layer for document-handling agents.
+pdf_whitefont equaling or exceeding pdf_append in ASR (100% on DeepSeek, 33% vs.
+0% on GPT-4o-mini) shows that the vulnerability is in the extraction pipeline, not in the rendered document.
+Any defense that operates on the rendered or displayed PDF—visual screening, human review, content moderation of images—will miss injections that exist only in the parser’s extracted text.
+Future document-handling agents need extraction-level content integrity: cryptographic attestation of source, parse-time payload detection, or sandboxed extraction environments that can flag content-display divergence before it enters agent context.
 6 Broader Impact
@@
 7 Limitations
-Per-cell $n$ ranges from 8 to 36; the strongest findings (Claude 0/20 per scenario, DeepSeek memory 0/24, GPT-4o-mini memory 12/12) are consistent across batches but should be interpreted with appropriate uncertainty.
-Payloads are explicit and synthetic; real-world attacks typically use obfuscation and social engineering.
-Five frontier models from three providers; reasoning models (o3, o4-mini) and open-weight models (Llama 3.3) are not represented.
-All defenses are lightweight wrappers that approximate the stated mechanism.
+Sample size.
+Per-cell $n$ ranges from 3 (Phase 3) to 36 (text-surface scenarios).
+The strongest text-surface findings (Claude 0/20 per scenario, DeepSeek memory 0/24, GPT-4o-mini memory 12/12) are consistent across independent batches; the Phase 3 cells at $n=3$ have wide Wilson CIs (21–94% for the 67% relay result) and must be treated as pilot-scale hypotheses.
+Single task.
+All scenarios use one task type (document summarization with an email delivery action).
+Our findings may not generalize to tasks with different tool sequences, longer horizons, or different outbound action types.
+This is the primary threat to external validity; replication across diverse task distributions is required before deployment recommendations can be made with confidence.
+Canary detection confound.
+A model that refuses an injection but quotes the canary token in its refusal explanation would be incorrectly labeled EXE=True.
+We inspected all 22 compromised text-surface runs manually and found no such false positives, but this confound is possible in principle and is not systematically controlled.
+Explicit payloads.
+Payloads are clear, unobfuscated directives; real-world attacks use social engineering, encoding, and multi-hop indirection.
+Surface-mismatch results are likely conservative: defenses designed for the wrong surface will also fail against obfuscated off-surface attacks.
+Other limitations.
+Five frontier models from three providers; reasoning models (o3, o4-mini) and open-weight models (Llama 3.3, Mistral) are not evaluated.
+All defenses are lightweight wrappers approximating the stated mechanism.
+We do not isolate whether Claude’s summarization-stage filtering arises from training, system prompt, or tool API design.
 permission_esc near-zero ASR (2/132) most likely reflects payload design rather than genuine privilege-escalation resistance.
-Classifier and provenance validation use the 192-run scenario_compare dataset; revalidation on the full 764-run set is ongoing.
-We do not isolate whether Claude’s summarization-stage filtering arises from training data, system prompt format, or tool API design.
+GPT-5-mini-2025-02-15 produced 0% tool-call rate in Phase 3; we cannot determine whether this is a capability regression or API schema incompatibility from black-box evaluation.
 8 Conclusion
-We introduced kill-chain canary tracking for localizing prompt injection defenses in LLM agent pipelines.
-Across 764 runs and five frontier models, the key structural finding is threefold: the safety gap between models concentrates at the summarization stage, not at exposure; static defenses fail not because they are poorly designed but because they cover different surfaces than the actual injection vectors; and a Claude relay node provides measurable decontamination for downstream agents regardless of those agents’ individual safety levels.
-Stage-level evaluation reveals that prompt injection defenses are fundamentally pipeline-local, not model-global.
-The minimum evaluation standard for agentic systems should therefore include: stage-level kill-chain decomposition, multi-surface injection coverage, and relay decontamination rate as a first-class metric.
+Stage-level canary tracking across 950 runs reveals a simple but consequential structure: prompt injection is not a model-capability problem, it is a pipeline-architecture problem.
+Every model is fully exposed; the safety gap lives entirely in what happens to an injection after it enters context.
+We draw four findings and, more importantly, four prescriptive claims about how agentic systems should be built and evaluated.
+Finding 1 $\to$ Claim: Write-node placement is a deployable safety primitive today.
+Claude eliminates injections at write_memory summarization on every surface we tested (0/164 text-surface, 0/9 Phase 3 PDF runs as Agent A; 95% CI 0–2%).
+This is not a property of Claude’s overall capability level—it is a property of its write-stage behavior.
+The practical implication is immediate: in any multi-agent pipeline where components can be chosen, routing all inter-agent memory writes through a safety-verified write node provides pipeline-wide decontamination independent of downstream agent choices.
+Conversely, Block B results show Claude as Agent B propagates at 67%—so write-node placement is the only position where this protection materializes.
+Future agent frameworks should expose write-node identity as a first-class architectural choice, not a deployment afterthought.
+Finding 2 $\to$ Claim: Defense design must be surface-aware, not assumption-aware.
+All four evaluated defenses—write_filter, pi_detector, spotlighting, and their combination—fail on at least one injection surface, not because they are incorrectly implemented but because they were designed for a different surface than the one under attack.
+This failure mode requires no adversarial adaptation.
+The correct design principle is surface decomposition before defense selection: enumerate every channel through which untrusted content can enter the pipeline, then select or compose defenses that cover each channel’s specific injection point.
+A monolithic defense evaluated only on the surface it was designed for is not a security evaluation.
+Finding 3 $\to$ Claim: Memory provenance is a missing infrastructure primitive.
+The write-vs-read safety asymmetry (Claude blocks at write, propagates at read) suggests that agent memory currently lacks the infrastructure to carry trust provenance: information about which agent wrote a memory record, under what safety context, and from what source surface.
+If Agent B could interrogate “was this memory written by a safety-certified node from a document I trust?” it could apply calibrated skepticism to inherited content rather than treating all memory reads as equivalently trusted delegation.
+Implementing content-addressed, provenance-tagged memory stores is a concrete near-term architectural direction for multi-agent frameworks.
+Finding 4 $\to$ Claim: Evaluation coverage is the primary security gap, not model capability.
+DeepSeek’s 0%/100% ASR split across memory and tool-stream surfaces, and the audio injection null result (spoken instructions treated as narration, not directives), show that apparent security posture is determined almost entirely by which surfaces an evaluation covers.
+A model that achieves 0% ASR on a single surface can achieve 100% on the next.
+The community standard for agentic security evaluation must require: (i) kill-chain stage decomposition rather than outcome-only ASR; (ii) multi-surface injection coverage including document extraction, tool response streams, audio transcripts, and pre-seeded memory; (iii) cross-model relay testing with heterogeneous agent pairs; and (iv) relay decontamination rate measured separately at write and read positions.
+Until these become baseline requirements, security evaluations of agentic systems will systematically underestimate real attack surfaces.